OpenStudy (anonymous):

A 1-m copper wire with a radius of 0.5 mm is stretched to a distance of 2 m. What is the fractional change in resistance ΔR/R as the wire is stretched?

OpenStudy (mos1635):

 V1=V2 (volume) πR1^2*l1=πR2^2*l2 R2=sqr(l1/l2)*R1=(1/4)*sqr2